Circular No. 7268

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999eb IN NGC 664 M. Modjaz and W. D. Li, University of California at Berkeley, on behalf

of the Lick Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a supernova in NGC 664. SN 1999eb was discovered on an unfiltered image taken on Oct. 2.4 UT (mag ∼ 16.2) and was confirmed on earlier images taken on Sept. 29.4 (mag ∼ 16.8), 22.4 (mag 18.0), and 19.4 (at the limiting mag of 19.0). The new object is located at α = 1^{h}43^{m}45^{s}.45, δ = +4^{o}13′25′′.9 (equinox 2000.0), which is 4′′.8 west and 4′′.1 north of the nucleus of NGC 664. A KAIT image of the same field on Sept. 13.4 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999eb. SN 1996bw and SN 1997W also appeared in NGC 664.

P. Garnavich, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum, taken by P. Berlind on Oct. 3.3 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows SN 1999eb to be a type-IIn supernova with narrow and intermediate-width Balmer emission lines superimposed on a blue continuum. He I 587.5-nm is also seen, but only as a broad component with a width of 4500 km/s (FWHM). The Balmer lines show a narrow absorption blueshifted by 900 km/s relative to the peak emission. Permitted Fe II features with P-Cyg profiles are also present. The spectrum is similar to SN 1995G (IAUC 6139).

SUPERNOVA 1999ec IN NGC 2207 Modjaz and Li also report the discovery of an apparent supernova in

NGC 2207 by the LOSS on unfiltered KAIT images taken on Oct. 2.5 and 3.5 (both with SN 1999ec at mag ∼ 17.9). SN 1999ec also appeared on an image taken on Sept. 20.5 (mag ∼ 17.5). The new object is located at α = 6^{h}16^{m}16^{s}.16, δ = −21^{o}22′09′′.8 (equinox 2000.0), which is 81′′.7 west and 12′′.1 north of the nucleus of NGC 2207. A KAIT image of the same field on 1998 Oct. 14.5 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999ec. SN 1975A also appeared in NGC 2207, which is a galaxy interacting with IC 2613.

SUPERNOVA 1999dj IN UGC 1258 Unfiltered CCD magnitude estimate by L. Kiss, K. Sárneczky, and G.

Szabó, Piszkéstető, Konkoly Observatory: Sept. 24.13 UT, 17.4 ± 0.3.

1999 October 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green